### v1

# On the parity of the Betti numbers of 3-manifolds with a parallel vector field

Abstract.
The question of whether a closed, orientable manifold can admit a nonzero vector field that is parallel with respect to some Riemannian metric is a classical problem in Differential Geometry, first posed by S. S. Chern [chern1966geometry]. In this work, we provide a complete answer to Chern’s question in dimension three. Specifically, we show that a closed, orientable 3-manifold admits a nonzero parallel vector field with respect to some Riemannian metric if and only if it is a Kähler mapping torus. Furthermore, we prove that the Betti numbers of any such 3-manifold are necessarily odd. A full classification of these manifolds is also obtained.
1. Introduction
In 1966, S. S. Chern [chern1966geometry] posed the problem of determining when a compact orientable manifold $M$ admits a nontrivial vector field that is parallel with respect to some Riemannian metric ([chern1966geometry]Example 3). He observed that the first two Betti numbers necessarily satisfy
$b_{1}(M)\geq 1,\qquad b_{2}(M)\geq b_{1}(M)-1,$
a condition now referred to as Chern’s criterion, and conjectured that these inequalities are not sufficient [chern1966geometry].
The situation for manifolds with boundary was clarified by P. Percell [percell1981parallel], who established that every compact manifold with boundary admits a vector field parallel with respect to some Riemannian metric. For closed manifolds, however, the problem remains considerably more intricate: identifying the precise topological conditions that are both necessary and sufficient for the existence of a nonvanishing parallel vector field is still a subtle challenge. It is known [karp1977parallel] that if a closed $n$-dimenional Riemannian manifold carries a parallel vector field, then its Euler–Poincaré characteristic must vanish. Furthermore, Bott [bott1967vector] demonstrated that all Pontryagin numbers of such a manifold are zero. Extending these observations, Karp [karp1977parallel] derived additional necessary conditions on the Betti numbers, namely
$b_{1}(M)\geq 1,\qquad b_{l+1}(M)\geq b_{l}(M)-b_{l-1}(M),\quad\text{for all }1\leq l\leq n-1.$
He also constructed an explicit example of a manifold satisfying both Chern’s criterion and Bott’s vanishing conditions on the Pontryagin numbers, yet admitting no parallel vector field for any Riemannian metric thus confirming Chern’s conjecture that the criterion is not sufficient.
A complete topological characterization was later obtained by Welsh [welsh1986manifolds], who showed that compact manifolds admitting a parallel vector field are precisely compact fiber bundles over tori with finite structural group.
The purpose of this note is to resolves Chern’s question in dimension $3$. We proves that the only $3$-dimensional closed orientable manifolds that admit a nonvanishing parallel vector field with respect to a Riemannian metric are the Kähler mapping torus. As a consequence, we show that their Betti numbers are necessarily odd, and we provide a full classification.
2. Basic Notions
2.1. Kähler Mapping Torus
We begin by recalling one of the fundamental constructions underlying our results.
Definition 2.1.
Let $\varphi\in\mathrm{Diff}(\Sigma)$ be a diffeomorphism of a closed, connected manifold $\Sigma$. The mapping torus of $\varphi$ is defined as
$\Sigma_{\varphi}=\frac{\Sigma\times[0,1]}{(x,0)\sim(\varphi(x),1)}.$
It forms the total space of a smooth fiber bundle
$\Sigma\hookrightarrow\Sigma_{\varphi}\xrightarrow{\;\pi\;}\mathbb{S}^{1}.$
If $(\Sigma,\Omega)$ is a symplectic manifold and $\varphi$ is a symplectomorphism, i.e. $\varphi^{*}\Omega=\Omega$, then $\Sigma_{\varphi}$ is called a symplectic mapping torus. Moreover, if $(\Sigma,\Omega)$ is a Kähler manifold endowed with a Hermitian structure $(J,h)$, and if $\varphi$ is a Hermitian isometry satisfying
$\varphi_{*}\circ J=J\circ\varphi_{*},\qquad\varphi^{*}h=h,$
(which automatically implies $\varphi^{*}\Omega=\Omega$), then $\Sigma_{\varphi}$ is referred to as a Kähler mapping torus [li2008topology].
3. Main Results
The main aim of this paper is to classify $M$ which carry a nontrivial vector field that is parallel with respect to a Riemannian metric.
Theorem 3.1.
Let $M$ be a closed, orientable, 3-manifold. Then $M$ admits a nontrivial parallel vector field with respect to some Riemannian metric if and only if $M$ is a Kähler mapping torus.
Proof.
(Necessary part) $(\Rightarrow)$: Let $V$ be a parallel vector field on $M$ with respect to a Riemannian metric $g$, and let $\alpha:=i_{V}g$ denote the associated $1$-form. Denote by $D$ the Levi-Civita connection of $g$. Since $DV=0$, it follows that $D\alpha=0$, and consequently $d\alpha=0$. Furthermore, the function $g(V,V)$ is constant on $M$; we denote this constant by $g(V,V)=c>0$.
Normalize $V$ and $\alpha$ by setting
$\widetilde{V}:=\frac{V}{\sqrt{c}},\qquad\widetilde{\alpha}:=\frac{\alpha}{\sqrt{c}}.$
Then $\widetilde{V}$ and $\widetilde{\alpha}$ satisfy
$\widetilde{\alpha}(\widetilde{V})=1,\qquad D\widetilde{V}=0,\qquad\mathcal{L}_{\widetilde{V}}g=0,\qquad g(\widetilde{V},\widetilde{V})=1.$
Next, consider the $2$-form
$\Omega:=i_{\widetilde{V}}v_{g},$
where $v_{g}$ denotes the Riemannian volume form associated with $g$. A straightforward computation gives
$d\Omega=d(i_{\widetilde{V}}v_{g})=\mathcal{L}_{\widetilde{V}}v_{g}=0.$
Moreover, we observe that
$i_{\widetilde{V}}(\widetilde{\alpha}\wedge v_{g})=0,$
which in turn implies
$v_{g}=\widetilde{\alpha}\wedge\Omega.$
Hence, $\widetilde{V}$ is the Reeb vector field of the cosymplectic structure $(\widetilde{\alpha},\Omega)$. Since $\widetilde{V}$ is Killing, it follows from [bazzoni2015k] that $(M,\widetilde{\alpha},\Omega)$ is a $K$-cosymplectic manifold. By [bazzoni2015k, Proposition 2.8], $M$ carries an almost contact metric structure $(\widetilde{\alpha},\phi,\widetilde{V},h)$ satisfying $L_{\widetilde{V}}h=0$. Furthermore, by [goldberg1969integrability, Proposition 3], this structure is normal ($N^{\phi}=0$), and hence $(M,\widetilde{\alpha},\widetilde{V},\phi,h)$ is a co-Kähler manifold. Finally, by [li2008topology, Theorem 2], $M$ is a Kähler mapping torus.
(Sufficient part) $(\Leftarrow)$: Consider a Kähler mapping torus $M=\Sigma_{\varphi}$. By [li2008topology, Theorem 2], $M$ admits a co-Kähler structure $(\phi,\xi,\eta,g)$ satisfying
$\phi^{2}=-I+\eta\otimes\xi,\qquad g(X,Y)=g(\phi X,\phi Y)+\eta(X)\eta(Y).$
Moreover, the Nijenhuis tensor vanishes:
$N^{\phi}(X,Y)=[\phi X,\phi Y]-\phi[\phi X,Y]-\phi[X,\phi Y]+\phi^{2}[X,Y]=0.$
As shown in [blair2010riemannian, blair1966theory, li2008topology, marrero1998new, cappelletti2013survey, olszak1981almost], every compact co-Kähler manifold satisfies
$\nabla^{LC}\eta=0,$
where $\nabla^{LC}$ denotes the Levi-Civita connection associated with $g$. Since $\eta(X)=g(X,\xi)$, a straightforward calculation then shows that the Reeb vector field $\xi$ is parallel:
$\nabla^{LC}\xi=0.$
This completes the first step of the proof. ∎
From Theorem 3.1, the only $3$-dimensional closed, orientable manifolds that admit a nonvanishing parallel vector field with respect to some Riemannian metric are the Kähler mapping torus. This completely answers Chern’s question in dimension $3$.
Yves Carrière [carriere1984flots] classified Riemannian flows on compact, orientable $3$-manifolds, and deduced that any such manifold admitting a nonvanishing Killing flow is a Seifert manifold. Now, consider a $3$-manifold $M$ endowed with a nonvanishing parallel vector field $V$ with respect to some Riemannian metric. Since $V$ is necessarily a Killing vector field, it follows from [carriere1984flots] that $M$ is a Seifert manifold. In what follows, we study the parity of the Betti numbers of $M$ and provide a complete classification of $3$-manifolds admitting a nonvanishing parallel vector field with respect to some Riemannian metric. This result may be viewed as a refinement of Carrière’s classification.
We can show the following result:
Theorem 3.2.
A closed, orientable $3$-manifold $M$ admitting a nontrivial parallel vector field with respect to some Riemannian metric has odd Betti numbers, and is diffeomorphic to one of the following types:
  1. (1)
    the product $\mathbb{S}^{2}\times\mathbb{S}^{1}$;
  2. (2)
    the three-dimensional torus $\mathbb{T}^{3}$;
  3. (3)
    one of the nontrivial torus bundles $\mathbb{T}^{2}_{A}$ over $\mathbb{S}^{1}$, where
    $A\in\left\{\begin{pmatrix}0&1\\ -1&0\end{pmatrix},\begin{pmatrix}-1&0\\ 0&-1\end{pmatrix},\begin{pmatrix}0&-1\\ 1&1\end{pmatrix},\begin{pmatrix}-1&-1\\ 1&0\end{pmatrix}\right\}$;
  4. (4)
    or a quotient of the form $(\mathbb{H}^{2}\times\mathbb{R})/\Gamma$, where $\Gamma\subset\mathrm{Isom}(\mathbb{H}^{2}\times\mathbb{R})$ is a discrete subgroup acting freely.
Proof.
By Theorem 3.1, a closed, orientable manifold $M$ admitting a nontrivial parallel vector field with respect to some Riemannian metric is a Kähler mapping torus. According to [li2008topology, Theorem 2], such a manifold $M$ is co-Kähler manifold. Furthermore, by [Chinea1993TopologyOC, Theorem 11], the first Betti number of $M$ is odd. Applying Poincaré duality, it then follows that all Betti numbers of $M$ are odd. This concludes the first step of the proof.
Since $M$ is a Kähler mapping torus, we have $M\simeq\Sigma_{\varphi}$, where $(\Sigma,J,h)$ is a Kähler manifold. Let $\mathrm{Isom}(\Sigma,h)$ denote the group of Hermitian isometries of $\Sigma$, and let $\mathrm{Isom}_{0}(\Sigma,h)$ denote its connected component of the identity. According to [bazzoni2014structure, Theorem 6.6], the diffeomorphism $\varphi$ is either isotopic to the identity or of finite order in the quotient group $\mathrm{Isom}(\Sigma,h)/\mathrm{Isom}_{0}(\Sigma,h)$; in particular, $\varphi$ is periodic. By [scott1983geometries, Theorem 5.4], $M$ is a Seifert bundle over a 2–orbifold $B$ with Euler number zero. Furthermore, as noted in [scott1983geometries], the three geometries $\mathbb{S}^{3}$, $\mathrm{Nil}$, and $\widetilde{\mathrm{SL}_{2}(\mathbb{R})}$ are excluded. Consequently, the classification of Seifert manifolds depends on the sign of the Euler characteristic $\chi(B)$. Case $\chi(B)>0$:
In this case, $M$ has $\mathbb{S}^{2}\times\mathbb{R}$ geometry. By [martelli2016introduction, Proposition 10.3.36], $M$ is diffeomorphic to one of $\mathbb{S}^{2}\times\mathbb{S}^{1}$, the orientable $\mathbb{S}^{1}$-bundle over $\mathbb{RP}^{2}$, or $\bigl(S^{2};(p,q),(p,-q)\bigr)$. All manifolds of the type $\bigl(S^{2};(p,q),(p,-q)\bigr)$ are diffeomorphic to $\mathbb{S}^{2}\times\mathbb{S}^{1}$. Moreover, by [martelli2016introduction, Exercise 10.3.37], the orientable $\mathbb{S}^{1}$-bundle over $\mathbb{RP}^{2}$ is diffeomorphic to $\mathbb{RP}^{3}\#\mathbb{RP}^{3}$, which has first Betti number $b_{1}(M)=0$ and is therefore excluded. Hence, the only possibility is $M\simeq\mathbb{S}^{2}\times\mathbb{S}^{1}$. Case $\chi(B)=0$: Here, $M$ has $\mathbb{R}^{3}$ geometry, and by [martelli2016introduction, Theorem 12.3.1], $M$ admits a flat metric. From the classification results of Bieberbach, Hantzsche–Wendt, and Wolf [hantzsche1935dreidimensionale, Wolf, bieberbach1911bewegungsgruppen, bieberbach1912bewegungsgruppen], the first homology group $H_{1}(M,\mathbb{Z})$ is isomorphic to one of
$\mathbb{Z}^{3},\quad\mathbb{Z}\oplus\mathbb{Z}_{2}^{2},\quad\mathbb{Z}\oplus\mathbb{Z}_{3},\quad\mathbb{Z}\oplus\mathbb{Z}_{2},\quad\mathbb{Z},\quad\mathbb{Z}_{4}^{2}.$
Among these six flat manifolds, the only one with $b_{1}(M)=0$ is the Hantzsche–Wendt manifold, for which $H_{1}(M,\mathbb{Z})\cong\mathbb{Z}_{4}^{2}$; it is therefore excluded. All remaining manifolds have a positive first Betti number, and by [marrero1998new, Theorem 3.2], $M$ is diffeomorphic to one of
$M_{1}(1),\ M_{2}(1),\ M^{\prime}_{1}(1),\ M^{\prime}_{2}(1),\ \text{or }\mathbb{T}^{3},$
in the notation of [marrero1998new]. Moreover, as shown in the examples of [marrero1998new], the remaining manifolds are torus bundles, explicitly given by
$M_{1}(1)$ $\simeq\mathbb{T}^{2}_{A},$ $A$ $=\begin{pmatrix}0&1\\ -1&0\end{pmatrix}$,
$M_{2}(1)$ $\simeq\mathbb{T}^{2}_{A},$ $A$ $=\begin{pmatrix}-1&0\\ 0&-1\end{pmatrix}$,
$M^{\prime}_{1}(1)$ $\simeq\mathbb{T}^{2}_{A},$ $A$ $=\begin{pmatrix}0&-1\\ 1&1\end{pmatrix}$,
$M^{\prime}_{2}(1)$ $\simeq\mathbb{T}^{2}_{A},$ $A$ $=\begin{pmatrix}-1&-1\\ 1&0\end{pmatrix}$.
Case $\chi(B)<0$: In this case, $M$ has geometry $\mathbb{H}^{2}\times\mathbb{R}$, and we have
$M\simeq(\mathbb{H}^{2}\times\mathbb{R})/\Gamma,$
where $\Gamma\subset\mathrm{Isom}(\mathbb{H}^{2}\times\mathbb{R})$ is a discrete group acting freely [scott1983geometries].
This completes the proof.
Corollary 3.1.
Let $M$ be a closed, orientable 3-manifold admitting a nontrivial parallel vector field $V$ with respect to some Riemannian metric $g$. Let $\alpha$ be the $1$-form metrically equivalent to $V$, and denote by $\mathcal{L}=\ker(\alpha)$ the corresponding foliation. Assume that the Ricci curvature of $g$ is non-negative on $\mathcal{L}\times\mathcal{L}$ and that $\mathcal{L}$ contains a compact leaf. Then the first Betti number $b_{1}(M)$ is either $1$ or $3$. Moreover, $b_{1}(M)=3$ if and only if $M$ is, up to a finite cover, diffeomorphic to the $3$-torus $\mathbb{T}^{3}$.
Proof.
By assumption, the codimension-one foliation $\mathcal{L}$ contains a compact leaf $L$. Then, by [tondeur1997geometry, Corollary 3.31], every leaf of $\mathcal{L}$ is diffeomorphic to $L$. Since the Ricci curvature of $g$ is non-negative on $\mathcal{L}\times\mathcal{L}$, it follows from [larz2011global, Theorem 2.82] that the first Betti number satisfies
$1\leq b_{1}(M)\leq 3.$
Combining this with Theorem 3.2, we deduce that $b_{1}(M)$ is either $1$ or $3$. Finally, by [schliebner2015lorentzian, Theorem 1.1], the equality $b_{1}(M)=3$ holds if and only if $M$ is diffeomorphic to the $3$-torus $\mathbb{T}^{3}$. ∎